OpenStudy (anonymous):

Giving medal for anwser! help ;p?? Find four solutions of the equation. Write the solutions as ordered pairs. 3x-y=-6 a. {(5,24), (-2,0),(-9,-18),(6,24)} b. {(5,21),(6,0),(-9,-21),(-2,24)} c. {(0,21),(-2,0),(-9,24),(5,24)} d. {(5,21),(-2,0),(-9,-21),(6,24)}

OpenStudy (whpalmer4):

Just try them out. For example, the first one in the first answer: 3(5)-24=-6 15-24=-6 -9=-6 That isn't true, so the fist answer is not the correct one. A correct answer will work for all of the listed ordered pairs.